OpenStudy (anonymous):

Composition of inverse functions f(x) = x^2+5, x<=0 g(x)=-sqrt(x-5), x>=5 for f(g(x)) I get x but for g(f(x)) I get -x I guess this makes sense looking at the domain and range of f(x), but it bothers me that I'm getting two different values(x and -x).

OpenStudy (zarkon):

\[g(f(x))=-\sqrt{f(x)-5}=-\sqrt{x^2+5-5}=-\sqrt{x^2}=-|x|\] since \(x\le 0\) we know that x is negative (or zero) so |x|=-x hence \(-|x|=-(-x)=x\)